OpenStudy (anonymous):

Do It Best Hardware is offering a 140-piece mechanic’s tool set plus a $65 tool chest for $226. What is the cost per tool?

OpenStudy (nurali):

cost of 140 tool and tool chest = 226 140x + 65 = 226 140x =226-65 140x =161 x=161/140 x = $1.15